8. f(x)=5x+4 and g(x)=7x-1
so I need to solve for this. (f+g)(x)

Found 2 solutions by ikleyn, fractalier:
Answer by ikleyn(52797)

.
Simply add these two binomials.


Answer by fractalier(6550)

(f+g)(x) = f(x) + g(x) = (5x+4) + (7x-1) = 12x + 3